# shapes,patterns,positioning

### 16 Feb 2024 v1

@@ 1 Introduction
In this paper we make several contributions to the development of simplicial homotopy theory, in particular the Kan-Quillen model structure, in an explicit or constructive style. Our starting point is the notion of an effective Kan fibration, which was introduced in BF (22): effective Kan fibrations are Kan fibration equipped with explicit solutions to lifting problems against horns satisfying a certain compatibility condition. The hope is that for these effective Kan fibrations we can obtain constructive and explicit versions of standard results in simplicial homotopy theory. In this paper we will show that that is indeed the case for two standard results.
@@
The second result is that Kan fibrations are the right class in a weak factorisation system. By now there is a well-established explicit version of a weak factorisation system, which goes by the name of an “algebraic weak factorisation system” (Gar (08); BG (16)); for our purposes the formulation of this notion as a “lifting algebraic weak factorisation system” Bou (23) is the most useful. In Section 4 we will show that the effective Kan fibrations are the right class in a lifting algebraic weak factorisation system.
@@ 2.2 Degenerate-preferring Kan fibrations
In this subsection, we shall introduce the degenerate-preferring Kan fibrations, but first we shall discuss the history behind the notion.
Kan fibrations were used in KL (21) to model dependent types for homotopy type theory. An essential part of this proof is that Kan fibrations are closed under pushforwards, so we can model dependent products. Homotopy type theory in turn models constructive mathematics. However, in BCP (15), it has been shown that a proof that Kan fibrations are closed under pushforwards cannot be made constructively. This problem has been addressed by moving away from simplicial sets, and instead use cubical sets, as in CCHM (18) and BCH (14). However, as simplicial sets are well-studied, it is still desirable to have a constructive notion of Kan fibrations in simplicial sets.
A first attempt to make the definition of Kan fibrations more constructive, would be to require functional choice of fillers. However, in Par (18), it was shown that these are also not constructively closed under pushforwards.
@@
At this point in time, there are two notions of Kan fibrations in simplicial sets which can constructively be shown to be closed under pushforwards, namely the effective Kan fibrations from BF (22) and the uniform Kan fibrations from GS (17). We will not discuss uniform Kan fibration here, but let us note that every effective Kan fibration is also a uniform Kan fibration, as shown in Corollary 11.2 of BF (22).
The effective Kan fibrations in turn are a superclass of the degenerate-preferring Kan fibrations, whose definition can be motivated by the following result of Van den Berg, Faber and Sattler. In Appendix C of BF (22) they show that if the lifting problem in Equation 7 has a degenerate solution, it is unique:
@@
Evan Cavello pointed out that degenerate-preferring Kan fibrations are similar to the “regular” or “normal” Kan fibrations in cubical sets (see for example the introduction of Swa (18)). Thierry Coquand based this name on the regular fiber spaces, which first appeared in Hur (55).
@@ 3.3 Some consequences
4 Lifting algebraic weak factorisation systems
The classical Kan fibrations are part of the “right class” in the weak factorisation system of the Kan-Quillen model structure. In this section, we will discuss a constructive variation on weak factorisation systems, namely lifting algebraic weak factorisation systems (lawfs). We will show that the symmetric effective and the effective Kan fibrations both fit into their own lawfs. In the first two subsections, we will recall the definition of a double category and a lawfs. These sections are based on Bou (23) and contain no original work. We will then define a class of double categories based on horn inclusions and study lawfss where the left double category comes from this class. We will then characterize the vertical morphisms of the right double category of such an lawfs and show that the effective and symmetric effective Kan fibrations satisfy this characterization.
@@ 4.1 Double categories
4.2 Lifting algebraic weak factorisation systems
@@
sending vertical morphisms $f$ to $(f,\phi_{-,f})$ and $(f,\phi_{f,-})$, respectively. The lifting structure $(\mathbb{L},\phi,\mathbb{R})$ on $\mathcal{C}$ is called a lifting algebraic weak factorisation system, or lawfs, if it satisfies the following two axioms:
@@
Let $\mathcal{C}$ be a locally presentable category and let $\mathbb{L}$ be a small double category over $\mathbb{S}\mathbf{q}(\mathcal{C})$. Then the canonical lifting structure $({\bf LLP}({\bf RLP}(\mathbb{L})),\phi,{\bf RLP}(\mathbb{L}))$ is a lifting awfs.
@@ 4.3 Degenerate-preferring Kan fibrations
In this section, we will construct a subcategory of $\mathbb{S}\textbf{q}(\textbf{sSet})$ which cofibrantly generates an lawfs where the vertical morphisms of the right double category are degenerate-preferring Kan fibrations.
@@ 4.4 The double category of horn pushout sequences
In the following sections, we will prove some general lemmas in order to construct an lawfs based on horn inclusions. To stay as general as possible, we will assume we are given a surjection $\ell:\mathcal{L}\to\Lambda$ be into the set of horn inclusions; in the cases of interest, we have $\mathcal{L}=\Lambda$ and $\ell=1_{\Lambda}$, or $\mathcal{L}$ the set of signed horn inclusions and $\ell=\pm$ the forgetful map which forgets the sign. In this section, we will build a double category $\mathbb{L}_{\ell}$ over $\mathbb{S}\textbf{q}(\textbf{sSet})$. We will later study right vertical maps of an lawfs cofibrantly generated by wide subcategories of $\mathbb{L}_{\ell}$.
@@
To make sure we include pullback squares, the horizontal morphisms in $\mathbb{L}_{\ell}$ between decidable sieves $S\subseteq\Delta^{a},T\subseteq\Delta^{b}$ are morphisms $f:[a]\to[b]$ in $\Delta$ such that $f^{*}(T)=S$. Horizontal compositions and identities are then taken in $\Delta$.
@@
Remark that for each $a\in\mathbb{N}$, there are only a finite amount of monomorphisms into $[a]$. Each of these correspond to a non-degenerate simplex of $\Delta^{a}$. The non-degenerate simplices that factor through a decidable sieve $S\subseteq\Delta^{a}$ can be counted. For a mono $f:S\hookrightarrow T$, we say that $f$ adds a non-degenerate simplex if it factors through $T$ but not through $S$.
@@
consists of a morphism $f:[a]\to[b]$ in $\Delta$ and a nondecreasing function $\mu:\mathbb{N}_{\leq l}\to\mathbb{N}_{\leq k}$ such that $\mu(0)=0$ and $\mu(l)=k$ such that for each $0\leq i\leq l$, we have that $f^{*}(T_{i})=S_{\mu(i)}$. For such squares, both horizontal domain and codomain correspond to $f$. Horizontal composition works by composing the sequences and the nondecreasing functions. Horizontal identities have vertical domain and codomain the empty horn pushout sequence and identity function on $\mathbb{N}_{\leq 0}$ as nondecreasing function.
@@ 5 Outlook
In general, we do not have a good understanding of the left double category of a cofibrantly generated lawfs, in particular, for the lawfs corresponding to (symmetric) effective Kan fibrations. One question is whether any vertical morphism in this left double category which corresponds to an inclusion sieves, is a horn pushout sequence. Another question in this direction is whether every pullback square of horn pushout sequences can be decomposed into face and degeneracy squares.

### 29 Feb 2024 v2

@@ 1 Introduction
In this paper we make several contributions to the development of simplicial homotopy theory, in particular the Kan-Quillen model structure, in an explicit or constructive style. Our motivation comes from homotopy type theory Uni (13), where Kan fibrations have been used in KL (21) to model dependent types for homotopy type theory. An essential part of this proof is that Kan fibrations are closed under pushforwards, so we can model dependent products.
Homotopy type theory in turn models constructive mathematics. However, in BCP (15) (see also Par (18)), it has been shown that a proof that Kan fibrations are closed under pushforwards cannot be made constructively. This problem has been addressed by moving away from simplicial sets, and instead use cubical sets, as in CCHM (18) and BCH (14). However, as simplicial sets are important and pervasive in homotopy theory, it is still desirable to have a constructive notion of Kan fibrations in simplicial sets.
At this point in time, there are two notions of Kan fibrations in simplicial sets which can constructively be shown to be closed under pushforwards, namely the effective Kan fibrations from BF (22) and the uniform Kan fibrations from GS (17). We will not discuss uniform Kan fibration here, but let us note that this notion of fibration is not “local”, as discussed in Appendix D of BF (22), and every effective Kan fibration is also a uniform Kan fibration, as shown in Corollary 11.2 of BF (22).
Therefore we will focus on the effective Kan fibrations, which are Kan fibration equipped with explicit solutions to lifting problems against horns satisfying a certain compatibility condition. The hope is that for these effective Kan fibrations we can obtain constructive and explicit versions of many standard results in simplicial homotopy theory, as was done in BF (22) for closure under pushforwards. In this paper we will show that that is indeed the case for two other standard results.
@@
The second result is that Kan fibrations are the right class in a weak factorisation system. By now there is a well-established explicit version of a weak factorisation system, which goes by the name of an “algebraic weak factorisation system” Gar (08); BG (16); Bou (23). In Section 4 we will show that the effective Kan fibrations are the right class in a algebraic weak factorisation system.
@@ 2.2 Degenerate-preferring Kan fibrations
In this subsection, we shall introduce the degenerate-preferring Kan fibrations, which are examples of functional Kan fibrations.
@@
The definition of the degenerate-preferring Kan fibrations can be motivated by the following result of Van den Berg, Faber and Sattler. In Appendix C of BF (22) they show that if the lifting problem in Equation 7 has a degenerate solution, it is unique:
@@
Evan Cavello has pointed out to us that degenerate-preferring Kan fibrations are similar to the “regular” or “normal” Kan fibrations in cubical sets (see for example the introduction of Swa (18)). Thierry Coquand based this name on the regular fiber spaces, which first appeared in Hur (55).
@@ 3.3 Some consequences
4 Algebraic weak factorisation systems
The classical Kan fibrations are part of the “right class” in the weak factorisation system of the Kan-Quillen model structure. In this section, we will discuss a constructive variation on weak factorisation systems, namely that of an algebraic weak factorisation system (awfs). We will show that the symmetric effective and the effective Kan fibrations both fit into their own awfs. In the first two subsections, we will recall the definition of a double category and an awfs; the definition of an awfs we give here is the one from Bou (23), where it is called a lifting algebraic weak factorisation system. Indeed, these sections are based on Bou (23) and contain no original work. We will then define a class of double categories based on horn inclusions and study awfss where the left double category comes from this class. We will then characterize the vertical morphisms of the right double category of such an awfs and show that the effective and symmetric effective Kan fibrations satisfy this characterization.
@@ 4.1 Double categories
4.2 Algebraic weak factorisation systems
@@
sending vertical morphisms $f$ to $(f,\phi_{-,f})$ and $(f,\phi_{f,-})$, respectively. The lifting structure $(\mathbb{L},\phi,\mathbb{R})$ on $\mathcal{C}$ is called an algebraic weak factorisation system, or awfs, if it satisfies the following two axioms:
@@
Let $\mathcal{C}$ be a locally presentable category and let $\mathbb{L}$ be a small double category over $\mathbb{S}\mathbf{q}(\mathcal{C})$. Then the canonical lifting structure $({\bf LLP}({\bf RLP}(\mathbb{L})),\phi,{\bf RLP}(\mathbb{L}))$ is an awfs.
@@ 4.3 Degenerate-preferring Kan fibrations
In this section, we will construct a subcategory of $\mathbb{S}\textbf{q}(\textbf{sSet})$ which cofibrantly generates an awfs where the vertical morphisms of the right double category are degenerate-preferring Kan fibrations.
@@ 4.4 The double category of horn pushout sequences
In the following sections, we will prove some general lemmas in order to construct an awfs based on horn inclusions. To stay as general as possible, we will assume we are given a surjection $\ell:\mathcal{L}\to\Lambda$ onto the set of horn inclusions; in the cases of interest, we have $\mathcal{L}=\Lambda$ and $\ell=1_{\Lambda}$, or $\mathcal{L}$ the set of signed horn inclusions and $\ell=\pm$ the forgetful map which forgets the sign. In this section, we will build a double category $\mathbb{L}_{\ell}$ over $\mathbb{S}\textbf{q}(\textbf{sSet})$. We will later study right vertical maps of an awfs cofibrantly generated by wide subcategories of $\mathbb{L}_{\ell}$.
@@
The horizontal morphisms in $\mathbb{L}_{\ell}$ between decidable sieves $S\subseteq\Delta^{a},T\subseteq\Delta^{b}$ are morphisms $f:[a]\to[b]$ in $\Delta$ such that $f^{*}(T)=S$; in other words, they are maps $f:[a]\to[b]$ fitting into pullback squares
[…]
Horizontal compositions and identities are then taken in $\Delta$.
@@
We remark that for each $a\in\mathbb{N}$, there are only a finite number of monomorphisms into $[a]$. Each of these correspond to a non-degenerate simplex of $\Delta^{a}$. The non-degenerate simplices that factor through a decidable sieve $S\subseteq\Delta^{a}$ can be counted. For a mono $f:S\hookrightarrow T$, we say that $f$ adds a non-degenerate simplex if it factors through $T$ but not through $S$.
@@
consists of a morphism $f:[a]\to[b]$ in $\Delta$ and a nondecreasing function $\mu:\mathbb{N}_{\leq l}\to\mathbb{N}_{\leq k}$ with $\mu(0)=0$ and $\mu(l)=k$ such that for each $0\leq i\leq l$, we have that $f^{*}(T_{i})=S_{\mu(i)}$.
For such squares, both horizontal domain and codomain correspond to $f$. Horizontal composition works by composing the sequences and the nondecreasing functions. Horizontal identities have vertical domain and codomain the empty horn pushout sequence and identity function on $\mathbb{N}_{\leq 0}$ as nondecreasing function.
@@ 5 Outlook
In general, we do not have a good understanding of the left double category of a cofibrantly generated awfs, in particular, for the awfs corresponding to (symmetric) effective Kan fibrations. One question is whether any vertical morphism in this left double category which corresponds to an inclusion of sieves, is a horn pushout sequence. Another question in this direction is whether every pullback square of horn pushout sequences can be decomposed into face and degeneracy squares.